# Local wind speed forecasting at short time horizons relying on both Numerical Weather Prediction and observations from surrounding stations

### v1

@@ 4.2.1 Fine tuning the global model for each station
As noted in the Introduction, Corsica is a Mediterranean island characterized by a highly heterogeneous, complex, and mountainous terrain, subject to diverse weather conditions. These factors contribute to significant spatial variability in wind regimes. Consequently, the performance of the model (e.g., ${\mathcal{M}}^{\prime}$), which has been globally trained using all stations, is expected to vary significantly across sites. To address this challenge and enhance model performance, we explore the possibility of fine-tuning its parameters for each station according to the methodology described in Sec. 3-3.4.
@@
Ajaccio $1$ $1.01$ $0.95$ $5.9$
$3$ $1.15$ $1.00$ $13.0$
$6$ $1.26$ $1.07$ $15.1$
Lucciana $1$ $1.08$ $1.03$ $4.6$
$3$ $1.24$ $1.16$ $6.5$
$6$ $1.35$ $1.22$ $9.6$
Figari $1$ $1.21$ $1.15$ $5.0$
$3$ $1.41$ $1.27$ $9.9$
$6$ $1.56$ $1.32$ $15.4$
Renno $1$ $0.97$ $0.92$ $5.2$
$3$ $1.10$ $1.00$ $9.1$
$6$ $1.14$ $1.05$ $7.5$
@@
Let us end this section with the results of Table 4 which compares the performance of the fine-tuned ${\mathcal{M}}^{\prime}$ model against a neural network model introduced in Baïle and Muzy (2023) and trained using the dataset described in Section 2-2.2, which spans a longer period than MeteoNet (2011-2020). This model has an architecture which is similar to the ${\mathcal{M}}_{GS}$ model decribed in Sec. 3-3.2 and takes, as input data, the last few hours observations from the target station and its 15 neighboring stations. This baseline model will be referred to as $\cal N$. Due to its relatively simple structure, $\cal N$ is trained and optimized separately for each station. The reported RMSE values for both models are computed over the same evaluation set to ensure a fair comparison. The results show that, across all sites and forecast horizons, ${\mathcal{M}}^{\prime}$ consistently outperforms $\cal N$, despite being trained on a significantly shorter time period. Furthermore, as expected, the performance gap widens with increasing forecast horizons. This highlights the advantage of ${\mathcal{M}}^{\prime}$, which integrates numerical weather prediction (NWP) data, over $\cal N$, which relies solely on past observed data. It also shows the interest of fine tuning using a shorter period when one cannot train a local model using a long enough period.
4.2.2 Predicting strong wind occurrences
Figure 10: Wind speed distributions observed at the Ajaccio (blue) and Bonifacio (orange) ground stations. Solid lines represent the corresponding Weibull distribution fits. Dashed lines indicate the 90 % quantiles for both locations.
One possible appealing application of our approach, which is both local and probabilistic, is to predict the occurrence of local strong wind regimes. Indeed, the notion of strong wind velocity is very dependent on the site. In Fig. 10, the observed wind velocity distributions in Ajaccio and Bonifacio sites are illustrated. We see that the wind distribution in Ajaccio is concentrated around low wind speeds, whereas in Bonifacio, the distribution is broader, indicating a higher likelihood of strong wind events. This reflects the distinct wind climates of the two locations: Bonifacio experiences significantly stronger and more turbulent winds. This is primary due to its exposure in the Strait of Bonifacio that channels and accelerates winds, particularly when they are northwesterly like Mistral. The high limestone cliffs and surrounding hills also induce sudden wind gusts, direction shifts, making the wind highly variable near the coast. In contrast, Ajaccio is more sheltered within its gulf, resulting in generally milder winds, with only occasional strong gusts from dominant regional wind patterns (notably during Mistral and Libeccio episodes in autumn and winter). The concept of “strong winds” can therefore vary significantly depending on the location. A practical approach to defining strong winds is to use a threshold based on a quantile of the specific wind speed distribution of each site $s$. This approach relates to an estimate of the potential wind impacts accounting for the adaptation of local infrastructure (Klawa and Ulbrich, 2003). The dashed lines in Fig. 10 represent the 90 % quantile values for both locations, meaning that only 10 % of wind occurrences exceed these thresholds at their respective sites.
Within this context, predicting strong wind events at a given location amounts to forecasting the occurrence of threshold exceedances, where the threshold corresponds to a high quantile relative to the site’s wind climatology. Let us denote $Q^{s}(\rho_{C},q)$ the quantile of probability level $q$ associated with the climatological distribution $\rho_{C}(V)$ relative to station $s$, i.e., the value such that:
[…]
Let ${\cal H}$ be the Heaviside function and define the function:
${\mathcal{H}}_{s,q}(z)={\mathcal{H}}\Big{(}z-Q^{s}(\rho_{C},q)\Big{)}=\begin{cases}1&\text{if }z\geq Q^{s}(\rho_{C},q),\\ 0&\text{otherwise}.\end{cases}$ (15)
For any site $s$, at time $t$ on day $d$, we consider the binary variable $I^{s}_{d,t}$ defined as
$I^{s}_{d,t}={\mathcal{H}}_{s,q}(V^{s}_{d,t})\;.$ (16)
This variable indicates the occurrence of strong wind speeds, characterized by exceeding a given threshold in the tail of the site’s climatological distribution. Predicting strong events occurrence thus amounts to predict the value of $I^{s}_{d,t+h}$ at each time $t$, day $d$, for each station $s$ and horizon $h$.
In order to assess the quality of the prediction of $I^{s}_{d,t}$, we use PSS defined in Eq. (12). One of its main advantages is that it is unaffected by class imbalance, making it particularly useful in forecasting of strong wind episodes.
Figure 11: Predicting strong event occurrences: Symbols ($\blacksquare$) represent deterministic predictions (based on mean value) while symbols ($\bullet$) represent probabilistic based prediction (based on quantile value). For each type of prediction, 4 curves are represented corresponding to a probability level $q=0.8,0.9,0.95,0.99$ that is used to define $I_{d,t}^{s}$ according to Eq. (16).
[…]
Figure 12: (a) Predicting strong event occurrences: comparisons of three models for predicting the exceeedance of 95 % quantile of each site. ()($\bullet$)( ∙ ) symbols represent the fine-tuned ${\mathcal{M}}^{\prime}$ model with predictor (18), ()($\blacksquare$)( ■ ) symbols represent the fine-tuned ${\mathcal{M}}^{\prime}$ model with predictor (17), () symbols represent the prediction from ${\mathcal{M}}_{AR}$ model and the ($\blacklozenge$) represent the performance of the linear prediction, where a linear regression over past observed wind speeds at the station and its surrounding stations is optimized, individually for each station, in order to minimize MSE. (b) The same model are compared in terms of CSI where the level $q=0.33$ is used in (18)).
In the case of a single point prediction at site $s$, day $d$ and time $t$, ${\widehat{{\bf{V}}}}_{d,t}^{s}$ as provided , for instance, by the ${\mathcal{M}}$ model in order to minimize the empirical MSE error or by the ${\mathcal{M}}^{\prime}$ model in order to minimize the estimated conditional MSE, a natural predictor of $I^{s}_{d,t+h}$ is
${\widehat{I}}^{s}_{d,t+h}={\mathcal{H}}_{s,q}({\widehat{V}}_{d,t+h}^{s})\;.$ (17)
On the other hand, for any station $s$, the model ${\mathcal{M}}^{\prime}$ provides, at each $t$ and for each horizon $h$, an estimate of ${\widehat{\rho}}^{s}_{d,t+h}$, the condition probability distribution of $V_{t+h}^{s}$ from which one can infer the conditional level-q quantile, $Q^{s}({\widehat{\rho}}^{s}_{d,t+h},q)$ as in Eq. (14). In this case, as shown in Mason (1979) (see also Jolliffe (2004)), one can explicitely build the “best” predictor of $I^{s}_{d,t+h}$, in the sense it is the one that maximizes the PSS, as:
${\widehat{I}}^{s}_{d,t+h}={\mathcal{H}}_{s,q}\Big{(}Q^{s}({\widehat{\rho}}^{s}_{d,t+h},q)\Big{)}\;.$ (18)
Such a prediction is different from the previous one in Eq. (17) since the expected conditional mean value ${\widehat{V}}^{s}_{t+h}$ has been replaced by the level-$q$ conditional quantile. In Fig. 11 we compare the PSS value, averaged for all 21 Corsican stations, as obtained on the test and validation periods, for the fine-tuned ${\mathcal{M}}^{\prime}$ model when one use estimators (17) and (18) for the extreme events occurrence. Comparison is performed for all horizons $h=1,2,3,4,5,6$ hours and threshold values corresponding to $q=0.8,0.9,0.95,0.99$ station climatology quantiles. The PSS performance is averaged across both the validation and test periods to provide more representative samples of extreme events, which becomes increasingly important as the quantile considered increases. We clearly observe that the method leveraging probabilistic prediction to estimate conditional quantiles, as in (18), achieves significantly better performance, particularly for extreme quantiles. As the threshold value $Q^{s}(\rho_{C},q)$ increases (i.e., as $q$ rises from 0.8 to 0.99), the predictive performance of the conditional mean ${\widehat{V}}_{t+h}^{s}$ deteriorates sharply, whereas the performance of the conditional quantile prediction $Q^{s}({\widehat{\rho}}_{t+h}^{s},q)$ improves. In other words, probabilistic forecasting effectively captures extreme wind speed occurrences, outperforming the standard mean squared error (MSE)-based predictor.
Let us finally assess the performances of the finely tuned probabilistic model ${\mathcal{M}}^{\prime}$ to predict strong wind speeds by comparing it (i) to its ”deterministic” counterpart, that is the fine tuned model ${\mathcal{M}}^{\prime}$ using the mean value as predictor, (ii) to a linear model $\cal L$ (see Sec. 3-3.6) optimized for each station and (iii) to the model ${\mathcal{M}}_{AR}$ defined in Sec. 3. The results for all sites in Corsica over both validation and test intervals are presented in Fig. 12 in the case $q=0.95$. Performance measures are assessed using both PSS (Fig. 12(a)) and CSI metrics defined in 3-3.5 (Fig. 12(b)). In the latter case, as advocated in Jolliffe (2004), the optimal probability threshold in Eq. (18) is no longer $q=0.95$ as for PSS but has to satisfy $q=\frac{CSI}{1+CSI}$, which corresponds to $q\approx 0.33$ for an expected CSI around $0.5$. Regarding the PSS, one observes in Fig. 12(a) that the ability of all the deterministic baseline models to predict strong wind events is significantly lower than that of the ${\mathcal{M}}^{\prime}$ model with an optimal threshold value. Incidentally, we can see that the model $\cal L$ outperforms ${\mathcal{M}}_{AR}$ at short time horizons. However, its performance declines rapidly as the forecast horizon increases which contrasts with ${\mathcal{M}}^{\prime}$ and ${\mathcal{M}}_{AR}$ models that both maintain stable predictive skill over time. Similar results can be observed in Fig. 12(b) for the CSI, although the benefits of ${\mathcal{M}}^{\prime}$ are less significant, particularly at largest horizons. This maybe due to the fact that the chosen threshold $q=0.33$, the same for all stations, is suboptimal in that case.

### v2

@@ 5.2.1 Fine tuning the global model for each station
As noted in the Introduction, Corsica is a Mediterranean island characterized by a highly heterogeneous, complex, and mountainous terrain, subject to diverse weather conditions. These factors contribute to significant spatial variability in wind regimes. Consequently, the performance of the model (e.g., $\mathcal{M}$$^{\prime}$), which has been globally trained using all stations, is expected to vary significantly across sites. To address this challenge and enhance model performance, we explore the possibility of fine-tuning its parameters for each station according to the methodology described in Sec. 4.4. For the probabilistic model $\mathcal{M}$$^{\prime}$, fine-tuning is performed using only the data corresponding to the target station. The globally trained model is loaded to provide initial weights for the training phase, and the model is then trained using the same architecture and negative log-likelihood loss (Eq. (7)). This operation is computationally inexpensive and can be completed on a standard laptop within minutes.
[…]
Figure 11: Relative RMSE improvement over the French administrative départements included in the MeteoNet Southeast dataset. Results are averaged over all stations in each département) and over all horizons.
As a remark, the relevance of the fine-tuning procedure is not limited to the region of Corsica; in fact it can be applied to any subset of data from the original dataset. To demonstrate the general benefit of this approach, fine-tuning was tested over all French administrative départements included in the Southeast MeteoNet dataset, retaining only those stations for which data were of sufficient quality. The relative RMSE improvement, averaged over all the stations in the département and over all forecast horizons, is illustrated in Figure 11. Observed improvements range between approximately $4\%$ and $11\%$.
5.2.2 Seasonal Performance Analysis and Comparison to model $\cal N$
Figure 12: Monthly relative RMSE of the fine-tuned $\mathcal{M}^{\prime}$ model, averaged over 21 Corsican stations and all forecast horizons (1–6 hours). Higher values indicate reduced model accuracy.
It is relevant to assess whether model performance exhibits seasonal variability. To this end, we computed the monthly relative RMSE (denoted as RRMSE and defined as RMSE divided by mean wind speed) of the fine-tuned $\mathcal{M}^{\prime}$ model, averaged over all 21 Corsican stations and forecast horizons (1–6 hours). Figure 12 displays the results. It can be observed that the relative RMSE is highest during the extended winter October–March, indicating reduced model accuracy during these months. This period is characterized by more dynamic and variable wind regimes, which may pose greater challenges for prediction. In contrast, the lowest relative RMSE values are observed during June-August, suggesting better model performance under more stable summer conditions with notably well established breeze regimes.
@@ 5.2.2 Seasonal Performance Analysis and Comparison to model $\cal N$
Ajaccio 11 $1.01$ $0.95$ $5.9$
33 $1.15$ $1.00$ $13.0$
66 $1.26$ $1.07$ $15.1$
Lucciana 11 $1.08$ $1.03$ $4.6$
33 $1.24$ $1.16$ $6.5$
66 $1.35$ $1.22$ $9.6$
Figari 11 $1.21$ $1.15$ $5.0$
33 $1.41$ $1.27$ $9.9$
66 $1.56$ $1.32$ $15.4$
Renno 11 $0.97$ $0.92$ $5.2$
33 $1.10$ $1.00$ $9.1$
66 $1.14$ $1.05$ $7.5$
Finally, Table 4 compares the performance of the fine-tuned $\mathcal{M}$$^{\prime}$ against a neural network model introduced in Baïle and Muzy (2023) which was trained using a dataset spanning a longer period (2011-2020) than that of MeteoNet. This model has an architecture which is similar to the ${\mathcal{M}}_{GS}$ model described in Sec. 4.2 and takes, as input data, the last few hours observations from the target station and its 15 neighboring stations. This baseline model will be referred to as $\cal N$. Due to its relatively simple structure, $\cal N$ is trained and optimized separately for each station. The reported RMSE values for both models are computed over the same evaluation set to ensure a fair comparison. The results show that, across all sites and forecast horizons, $\mathcal{M}$$^{\prime}$ consistently outperforms $\cal N$, despite being trained on a significantly shorter time period. Furthermore, as expected, the performance gap tends to widen with increasing forecast horizons. This highlights the advantage of $\mathcal{M}$$^{\prime}$, which integrates numerical weather prediction (NWP) data, over $\cal N$, which relies solely on past observed data. It also shows the importance of fine-tuning, particularly when training is conducted on a dataset that is too limited to fully optimize the model for local conditions.
5.2.3 Predicting strong wind occurrences
Figure 13: Wind speed distributions observed at the Ajaccio (blue) and Bonifacio (orange) ground stations. Solid lines represent the corresponding Weibull distribution fits. Dashed lines indicate the 90 % quantiles for both locations. Wind speed units are $ms^{-1}$.
One possible appealing application of finely-tuned model $\mathcal{M}$$^{\prime}$, which is both local and probabilistic, is to predict the occurrence of local strong wind regimes. Indeed, the notion of strong wind velocity is very dependent on the site. In Fig. 13, the observed wind velocity distributions in Ajaccio and Bonifacio are illustrated. We see that the wind distribution in Ajaccio is concentrated around low wind speeds, whereas in Bonifacio, the distribution is broader, indicating a higher likelihood of strong wind events. This reflects the distinct wind climates of the two locations: Bonifacio experiences significantly stronger and more turbulent winds. This is primary due to its exposure in the Strait of Bonifacio that channels and accelerates winds, particularly when they are northwesterly like Mistral. The high limestone cliffs and surrounding hills also induce sudden wind gusts and direction shifts, making the wind highly variable near the coast. In contrast, Ajaccio is more sheltered within its gulf, resulting in generally milder winds, with only occasional strong gusts from dominant regional wind patterns (notably during Mistral and Libeccio episodes in autumn and winter). The concept of “strong winds” can therefore vary significantly depending on the location. As explained in Sec. 3.3, a practical approach to defining strong winds is to use a threshold based on the specific wind speed distribution of each site $s$. This approach relates to an estimate of the potential wind impacts accounting for the adaptation of local infrastructure (Klawa and Ulbrich, 2003). The dashed lines in Fig. 13 represent the 90 % quantile values for both locations, meaning that only 10 % of wind occurrences exceed these thresholds at their respective sites.
Within this context, predicting strong wind events at a given location amounts to forecasting the occurrence of threshold exceedances corresponding a high quantile relative to the site’s wind climatology. In Sec. 3.3, such a quantity is denoted as $I_{s,d,t}(Q_{C}^{s}(p))$ (Eq. (10)) with a probability level $p$ close to $p=1$.
In order to assess the quality of the prediction, we use PSS as discussed in Sec. 3.3 (Eq. (11)). One of its main advantages is that it is unaffected by class imbalance, making it particularly useful in forecasting of strong wind episodes.
Figure 14: Predicting strong event occurrences: Symbols ($\blacksquare$) represent deterministic predictions (based on mean value) while symbols ($\bullet$) represent probabilistic based prediction (based on quantile value). For each type of prediction, 4 curves are represented corresponding to a probability level $p=0.8,0.9,0.95,0.99$ that is used to define $I_{s,d,t}$ according to Eqs. (13) and (15).
@@ 5.2.3 Predicting strong wind occurrences
In the case of a single point prediction at site $s$, day $d$ and time $t$ the corresponding estimator is denoted as ${\widehat{I}}^{1}_{s,d,t+h}$ (Eq. (13)) while estimator ${\widehat{I}}^{2}_{s,d,t+h}$ (Eq. (15)) is built from a probabilistic output of model $\mathcal{M}$$^{\prime}$ in order to optimize PSS . In Fig. 14 we compare the PSS value, averaged for all 21 Corsican stations, as obtained on the test and validation periods, for the fine-tuned $\mathcal{M}$$^{\prime}$ model when one use estimators ${\widehat{I}}^{1}_{s,d,t+h}$ and ${\widehat{I}}^{2}_{s,d,t+h}$. Comparison is performed for all horizons $h=1,2,3,4,5,6$ hours and threshold values corresponding to probability levels $p=0.8,0.9,0.95,0.99$ station climatology quantiles. The PSS performance is averaged across both the validation and test periods to provide more representative samples of extreme events, which becomes increasingly important as the quantile considered increases. We clearly observe that $I^{2}_{s,d,t+h}$, the prediction obtained by leveraging probabilistic forecasting to estimate conditional quantiles, achieves significantly better performance, particularly for the extreme quantiles. As the threshold value $Q_{C}^{s}(p)$ increases when $p$ rises from 0.8 to 0.99, the predictive performance of ${\widehat{I}}^{1}_{s,d,t+h}$ (based on the conditional mean ${\widehat{V}}_{t+h}^{s}$, see Eq. (13)) deteriorates sharply, whereas the performance of ${\widehat{I}}^{2}_{s,d,t+h}$ (based on the conditional quantile prediction $\widehat{Q}^{s}_{t+h}(p)$, see Eqs (14) and (15)), improves. In other words, probabilistic forecasting effectively captures extreme wind speed occurrences, outperforming the standard mean squared error (MSE)-based predictor.
@@
6 Discussion about some limitations and specific issues.
6.1 Temporal and spatial limitations
It should be noted that the results of this study are based on a limited three-year temporal period (2016-2018). Such a period is clearly too short to capture all wind regimes, extreme occurrences or interannual variabilities. However, we can first emphasize that the objective in this study is to design and evaluate a short-term wind speed forecasting method, not to assess the long-term wind resource potential or climatological characteristics of specific sites. In contrast to wind resource assessment studies—where long historical datasets are essential to estimate inter-annual variability and long-term extremes—this work focuses on optimizing a predictive model that delivers accurate short-term probabilistic forecasts using available real-time information. From this perspective, the key concern is not the long-term frequency of extremes, but rather the model’s ability to generalize to a broad range of conditions within the training and evaluation window. Second, the proposed neural network architecture is trained to optimize full accuracy across the entire distribution of wind speeds. Notably in its probabilistic version that performs better for strong events, it leverages all available data to learn the parameters of the forecast distribution. Unlike models designed specifically to predict rare or extreme events—often relying on tail-focused objectives or imbalance-aware training strategies—the approach proposed in this work does not require over-representation of extremes in the training data. Instead, it is designed to perform well across the full spectrum of possible outcomes, including high wind speeds when they appear. Last, the extremes we consider in the evaluation are defined with respect to empirical quantiles computed from the 3-year dataset used in the study. That is, we use threshold exceedances defined within the climatology of the available period (e.g., 95th percentile of observed wind speeds at each site). By construction, such thresholds ensure that extreme events, as defined in this study, are present in the dataset and therefore adequately represented in the training and testing phases. This allows us to meaningfully assess the model’s performance under “extreme” conditions, even within a 3-year window. However, extending the evaluation to a longer historical period would help further validate the robustness of the method under broader climatological variability.
Regarding the spatial scope of this study, the results presented in Fig. 11 indicate that the RMSE gains obtained when fine-tuning the globally trained model are consistent across the various subregions of the southeastern MeteoNet domain, and comparable to those observed in Corsica. This outcome suggests that the proposed two-step training and fine-tuning strategy is transferable and can be successfully replicated within other parts of the same broad climatic zone. Nevertheless, we recognize that regional variability, particularly across distinct climatic regimes, may influence the model’s performance. Extending the evaluation to the northwestern (Atlantic) regions of MeteoNet, characterized by meteorological conditions markedly different from those of the Mediterranean, represents an important direction for future work.
6.2 Considerations for wind energy applications
In the context of wind energy forecasting, it is important to recognize that standard meteorological stations typically measure wind speed at heights around 10 m, whereas wind turbine hub heights are commonly between 80 and 140 m. This vertical discrepancy presents a known challenge in operational forecasting. While the models proposed in this paper are trained using near-surface observations and NWP fields, their relevance to wind energy applications is supported by established extrapolation techniques. These include empirical methods such as the power law and logarithmic wind profiles (Gualtieri, 2015; Lopez-Villalobos et al., 2022), as well as site-specific shear estimates from the integration, with long-term mast observations, of reanalysis data (Gualtieri, 2021; Yang et al., 2024) or short-term LiDAR measurements (Floors et al., 2011). Additionally machine learning approaches capable of capturing non-linear atmospheric dynamics (Gualtieri, 2019) have been successfully used to model vertical wind profiles and extend near-surface forecasts to hub height. While these methods enhance the practical utility of near-surface forecasts, further validation based on actual measurements at hub height remains necessary (Drechsel et al., 2012) and is identified as an interesting prospect for future work.
One can mention another limitation of the present study: although its evaluation focuses on point-wise accuracy metrics such as measured by RMSE, it is important to note that operational wind energy applications are highly sensitive not only to the mean forecast accuracy but also to the temporal structure and variability of wind speed. In this context, a detailed analysis of the spectral content and temporal correlation structure of the predicted wind signals is crucial for properly characterizing ramping behavior and short-term fluctuations. As a perspective, future work will be dedicated to computing variance-based metrics (e.g., standard deviation ratio, normalized RMSE) and conducting spectral error analyses, in order to assess how well the models $\mathcal{M}$ and $\mathcal{M}^{\prime}$ capture the multi-scale temporal variability of wind fluctuations.
6.3 Operational feasibility
Table 5: Summary of Hourly Prediction Chain (with 6-hour NWP refresh)
Step Frequency Description Execution Time
NWP forecast availability Every 6 h From meteorological center 1–3 h
GRIB file download Every 6 h Transfer NWP forecasts locally $\sim$1 min
Ground station data update Every hour Load observations from file or API $\sim$1 min
Feature preparation Every hour Normalize and assemble input tensors (NWP + station) $\sim$1 min
NN model inference Every hour Forecast wind speed for 1–6 h lead times $\sim$5 s
To assess the practical applicability of the wind speed forecasting approach developed above, one can outline a theoretical operational prediction chain aligned with current production cycles of NWP systems such as AROME and ARPEGE. These models typically provide forecasts four times daily (i.e., initialized at 00, 06, 12 and 18 UTC), which become available approximately 1–3 hours later. Once available, GRIB files are downloaded and parsed, and the relevant fields are stored locally for use throughout the subsequent 6-hour cycle.
In the framework proposed by this study, the hybrid model $\mathcal{M}$} or $\mathcal{M}$$^{\prime}$—designed to produce forecasts from 1 to 6 hours ahead—is executed hourly. At each prediction time, the latest ground station measurements are retrieved, and input tensors are assembled from the stored NWP fields and recent observations from nearby ground stations. The inference step is computationally lightweight: on a standard laptop, the entire prediction cycle—including data loading, preprocessing, and model inference—takes less than 5 seconds.
Table 5 summarizes the timing and sequence of operations in the hourly prediction loop. This architecture ensures low-latency and high-frequency forecasting, compatible with real-time operational constraints. By leveraging pre-stored NWP fields and hourly updates from ground stations, the system maintains responsiveness and computational efficiency, making it suitable for applications such as wind energy management, local warnings, and short-term response.